### v2

# EgoAgent: A Joint Predictive Agent Model in Egocentric Worlds

Abstract
This paper addresses the task of learning an agent model behaving like humans, which can jointly perceive, predict, and act in egocentric worlds. Previous methods usually train separate models for these three abilities, which prevents them from learning from each other. In this paper, we propose a joint predictive agent model, named EgoAgent, that simultaneously learns to represent the world, predict future states, and take reasonable actions within a single transformer. EgoAgent introduces two innovations to learn from the causal and temporally intertwined nature of these abilities: (1) Interleaved sequential modeling of states and actions with the causal attention mechanism, and (2) A joint embedding-action-prediction architecture featuring temporal asymmetric predictor-observer branches. Integrating these designs based on JEPA [29], EgoAgent unifies these capabilities in a cohesive learning framework. Comprehensive evaluations of EgoAgent on representative tasks such as image classification, egocentric future state prediction, and 3D human motion prediction tasks demonstrate the superiority of our method. The code and trained model will be released for reproducibility.
Figure 1: (a) Humans learn from the world by continuously perceiving the egocentric world, predicting future states, and taking actions to achieve their goals. (b) Inspired by the human learning process, we proposed EgoAgent, a joint predictive agent model in egocentric worlds that can learn to represent, predict, and act from egocentric video and 3D human skeleton sequences.
1 Introduction
Cognition is embodied; it arises from bodily interactions with the world,” as stated by Linda B. Smith [^1] Linda Brawn Smith is a cognitive scientist recognized for her work in developmental psychology and cognitive science, with over 52,000 citations. She won the David E. Rumelhart Prize for theoretical contributions to cognitive science and is a member of both the National Academy of Sciences and the American Academy of Arts and Science. in [43]. This perspective emphasizes that cognition emerges from continuous, dynamic interactions between the brain, body, and environment, rather than relying on symbolic systems detached from real-world experiences [42, 8, 39]. Through these ongoing interactions, humans naturally acquire foundational abilities such as visual perception, predictive understanding of world dynamics, and action anticipation. As highlighted by Yann LeCun [29], replicating these abilities in artificial systems remains one major challenge in AI research. Motivated by this human learning process, we propose an agent model that jointly learns these abilities from egocentric RGB observations and their responsive actions. Such abilities are essential for applications in robotics, augmented reality, and gaming where agents must understand what is happening in the world, anticipate what will happen next, and decide what actions to take.
Existing approaches typically address these capabilities in isolated tasks: (1) Visual representation learning [10, 26, 46], which encodes high-level representations of human observations of the world; (2) World models [36, 52], which learn predictive representations of world state transitions conditioned on actions. (3) Action prediction [35, 9], which forecasts future human actions based on prior motion. However, research in cognitive science, particularly the famous Common Coding Theory [38, 8], claims that perception and action are not separate processes but deeply intertwined and mutually reinforcing, operating within a shared representational space. Therefore, modeling these abilities into a unified framework can be not only mutually beneficial but also fundamentally feasible for agent models to achieve human-like understanding in egocentric worlds.
While combining these abilities offers clear benefits, integrating them into one model remains challenging. Humans interact with the world through an interleaved cycle of perceiving egocentric observations and taking actions, creating close interdependencies between observations and actions in time and causality. To mimic and learn from this behavior, researchers must model the causal and temporally interlaced nature of observations and actions within the network, and design effective supervision signals to capture their dependencies, which is an unsolved problem.
In response, we propose EgoAgent, a joint predictive agent model that simultaneously learns to represent world observations, predict future states, and act based on learned representations (Fig. 1). To model and learn from the intertwined nature of these abilities, EgoAgent introduces two key innovations: (1) Formulating human-world interactions as interleaved sequences of “state-action-state-action” tokens processed by the causal attention mechanism. This design explicitly captures the interlaced structure of egocentric observations and responsive actions across temporal and causal dimensions within a unified representational space. (2) A joint embedding-action-prediction (JEAP) architecture with temporal asymmetric predictor-observer branches, where the predictor branch leverages past information to predict future world states while the observer extracts the target world states from raw observations. By minimizing the prediction error in feature spaces, JEAP is tailored to learn the temporal transitions of the world, thereby integrating perception, action, and prediction into a cohesive learning framework.
Extensive experiments show that EgoAgent effectively handles perception, prediction, and action tasks, even surpassing state-of-the-art methods across individual benchmarks. Specifically, it outperforms the leading egocentric video pretraining method [44] by +1.32% Top1 accuracy on ImageNet-1K [15] for image classification and +3.90% success rate of moving cube in the TriFinger simulator [48]. [^2] Image classification serves as a standard benchmark for visual perception, while TriFinger evaluates the transferability of learned representations to egocentric manipulation with a three-finger robot. Together, these tasks assess a model’s ability to learn meaningful visual representations. On Ego-Exo4D [20], it achieves +16.28% Top1 accuracy and +16.95% mAP for future state prediction, while also exceeding 3D human motion prediction methods by improving MPJPE by -0.82 at a 30 fps prediction rate.
Our contributions are as follows: (1) We develop the first agent model that can simultaneously learn to represent egocentric observations, predict future states, and generate actions in a unified representational space. (2) We propose a JEAP architecture with temporal asymmetric branches to learn both semantic understanding and temporal transitions of the dynamic multi-modal egocentric world. (3) We evaluate our EgoAgent on egocentric video-action datasets and show its superior abilities in visual representation, world states prediction, and 3D human motion prediction tasks.
2 Related Work
Figure 2: The overall framework of EgoAgent. EgoAgent adopts the Joint Embedding-Action-Prediction Architecture to project input egocentric video frames $I$ and 3D human poses $A$ into image tokens $i$ and action tokens $a$. Learnable action query token $q_{a}$ and future state query token $q_{s}$ are appended after the image and action tokens to stimulate EgoAgent’s predictor $\Theta_{pred}$ to predict the next action $A^{\prime}$ and world state $S^{\prime}$. The action prediction loss $\mathcal{L}_{act}$ is used to supervise the predicted actions. Since EgoAgent predicts world states $S^{\prime}$ in a semantic feature space, the target world state $S$ is obtained from a momentum observer $\Theta_{obs}$. Features from the predictor and observer branches are aligned through a state prediction loss $\mathcal{L}_{pred}$. When 3D poses are not paired with video frames, EgoAgent randomly crops different views from the video and learns representative features only with image tokens $i$ and state query tokens $q_{s}$ by a representation loss $\mathcal{L}_{rep}$. For easier understanding, we only illustrate the supervisions on predicting the scene state and action at time step $t$.
World Models and Agent Models. Ha and Schmidhuber [24] defined the world model as a predictive representation of the environment, while the agent model utilizes this learned representation to make decisions and take actions. Recently, generative world models [52, 2, 49] have gained significant attention, often employing an autoencoder framework to simulate transitions between video frames. These models have demonstrated success in applications such as autonomous driving [47], robotics [23], and game control [5]. For instance, GAIA-1 [27] maps multi-modal driving signals into discrete tokens, and processes them using an autoregressive transformer for unified representation learning. In contrast, predictive world models [3, 7, 18] focus on learning state transitions through the joint embedding predictive architecture (JEPA) [29] without generating videos explicitly. For example, MC-JEPA [6] incorporates self-supervised learning with optical flow estimation to learn content features and motion dynamics. In this work, we extend the JEPA world model into an agent model in egocentric worlds, which not only predicts world states based on human actions but also predicts actions based on learned world state transitions.
Egocentric Visual Representation Learning. Traditional visual representation methods learn the features on large-scale curated image datasets, such as ImageNet [15], using either supervised [40, 28, 46] or self-supervised methods [25, 10, 12]. However, these datasets primarily consist of third-person perspectives, which do not fully align with how humans perceive the world. To bridge this gap, researchers have turned to egocentric videos, which offer a more natural and immersive viewpoint for learning visual representations. Models like R3M [37] and VIP [31] leverage the temporal relationships in egocentric videos to learn generalizable representations from Ego4D [19] for robotics applications. Recently, DoRA [44] learns object-level representations by tracking movements on an egocentric video dataset, WalkingTours. In this paper, we propose an agent model that not only learns robust representations from egocentric videos but also predicts future states and actions.
3D Human Motion Prediction. Human motion prediction models [30, 50, 45] aim to forecast future body movements given historical poses. Aksan et al. [1] proposed a transformer-based architecture to capture the spatial-temporal dependencies of human motion over short and long horizons. Cui et al. [14] represent the human skeleton as a dynamic graph to adaptively learn joint connection strengths for better prediction accuracy. Other works [32, 53] incorporate physics-based priors to ensure the generated motions adhere to physical principles. Recent research has integrated multimodal cues such as textual instructions [34], eye gaze [54], and 3D objects [51] as conditions for human motion prediction. Notably, Cao et al. [9] proposed a three-stage framework that consists of goal prediction, path planning, and pose finalization, which effectively enhances scene-aware motion prediction.
3 Method
Our goal is to develop an agent model that can represent world states, predict future states, and generate the following actions by simulating how humans interact with the world through egocentric vision and 3D skeletal motion. The overview of our framework is presented in Fig. 2. We first introduce our joint embedding-action-prediction architecture in Sec. 3.1. We then introduce how EgoAgent can be trained in Sec. 3.2. Finally, we demonstrate the application of EgoAgent in the three tasks in Sec. 3.3.
3.1 Joint Embedding-Action-Prediction Architecture for EgoAgent
Given egocentric observations $V=\{I_{1},I_{2},...I_{t}\}$ and history human actions $A=\{A_{1},A_{2},...A_{t}\}$, our agent model aims to predict the future world state $S_{t+1}$ associated with the next observation $I_{t+1}$ while simultaneously generate plausible human action $A_{t+1}$ by understanding world states $S_{[0:t]}$. The world states are encoded as semantic embeddings, whereas human actions are represented as 3D poses.
Joint prediction pipeline. Our pipeline begins by encoding the egocentric video $V$ and human action sequence $A$ into semantic features. For egocentric video, we follow the approach of ViT [16], applying a convolutional layer to each video frame $I_{t}$ to produce a feature map. This map is then subdivided and flattened to obtain the final image feature $i_{t}$. For human actions, we utilize a convolutional layer followed by layer normalization (LN) and a Gaussian error linear unit (GeLU) activation to map the 3D skeletons $A_{t}$ into action features $a_{t}$. These feature vectors are then processed by a causal transformer network to generate the target outputs. As illustrated in Fig. 2, the input sequence is structured as a sequence of image, action, and query tokens.
At each time step $t$, we assemble a structured sequence of tokens, including the image token $i_{t}$, an action query token $q_{a}$, the action token $a_{t}$, and a future state query token $q_{s}$, where the query tokens are learnable embeddings. Enabled by the causal attention mechanism of the transformer, $q_{s}$ prompts the model to incorporate both previous and current image tokens $i_{[0:t]}$ and action tokens $a_{[0:t]}$ for predicting $s_{t+1}$. Notably, we insert $q_{a}$ right after the image tokens and before the action tokens to prevent the model from accessing the future action $a_{t}$ during prediction. This ensures that the model relies solely on past observations $i_{[0:t]}$ and past actions $a_{[0:t-1]}$ when generating future actions. Subsequently, the transformer network processes these tokens and outputs the action embedding $a_{t}^{{}^{\prime}}$ and future state embedding $s_{t+1}^{{}^{\prime}}$ in response to the query tokens. Finally, we utilize separate MLP networks to map these embeddings to the predicted actions $A_{t}^{{}^{\prime}}$ and future world states $S_{t+1}^{{}^{\prime}}$.
JEPA with temporal asymmetric branches. Unlike actions, which can be directly obtained and supervised, world states must be learned from observations while also be used as supervision signals for state prediction. To achieve this, we propose a temporal asymmetric predictor-observer architecture based on JEPA [29, 18], a framework that combines representation learning with predictive learning. Despite of the predictor branch executing the joint prediction pipeline, we introduce an observer branch that only processes image inputs by extracting current-frame features for representation learning supervision and providing next-frame features for future state supervision alongside the aforementioned predictor branch. Specifically, the observer branch takes image tokens $i_{t}$ and state query tokens $q_{s}$ as input, and outputs current state embeddings $s_{t}$, which is fed into the same MLP network used for image embeddings in the observer branch to obtain world states $S_{t}$.
Learning in semantic feature space. Notably, unlike previous methods [4, 52] that rely on pretrained reconstruction-based tokenizers such as VQGAN [17] to convert images into discrete tokens, we employ learnable convolutional layers to project images into continuous semantic embeddings. Discrete tokens prioritize pixel-level reconstruction over high-level semantic information, whereas humans make predictions based on abstract concepts rather than by picturing at the pixel level. Therefore, our approach aligns more closely with human reasoning, enhancing the model’s performance in future state prediction.
3.2 Training EgoAgent
Base model. We employ InternLM, an open-source LLM [41], as the foundational architecture for EgoAgent, adapting it to the JEAP architecture mentioned in Sec. 3.1. This is motivated by two key considerations: (1) Our objective of near-future world state and action prediction naturally aligns with the next-token prediction mechanism inherent in LLMs. (2) The structured yet flexible architecture of LLMs provides a scalable foundation for our specific tasks. We do NOT initialize the LLM with any pretrained weights, as we consider that visual perception and prediction abilities can be learned without language priors [4].
Supervisions for predicting world states and actions. EgoAgent consists of two branches: a predictor branch, which predicts future world states $S_{t+1}^{{}^{\prime}}$ and actions $A_{t}^{{}^{\prime}}$, and an asymmetric observer branch, which extracts the future world states $S_{t+1}$ using only the egocentric image $I_{t+1}$. Given the input token sequence at time step $t$, the loss functions for future state and action prediction are defined as:
$\small\mathcal{L}_{act}(t)$ $=\mathcal{L}_{1}(A_{t}^{\prime},A_{t}),$ (1)
$\mathcal{L}_{pred}(t)$ $=\mathcal{L}_{dino}(S_{t+1}^{\prime},sg[S_{t+1}]),$ (2)
where $sg[\cdot]$ denotes the stop-gradient operation, $\mathcal{L}_{1}$ and $\mathcal{L}_{dino}$ represent the L1 loss and the DINO loss [10], respectively. Following common practices in self-supervised learning [10, 25], sg[$\cdot$] is applied to block the gradients from back-propagating to the teacher branch. The weights of the observer branch are updated in each iteration using an exponential moving average (EMA) of the predictor branch.
Self-supervision for learning powerful representations. When humans learn to interact with the environment, they first develop an understanding of the observed world and objects, which then aids them in predicting future states of the world and making appropriate responses. Inspired by this process, we introduce an additional self-supervised learning loss on EgoAgent to facilitate learning representative features from egocentric videos from scratch:
$\small\mathcal{L}_{rep}(t)$ $=\mathcal{L}_{dino}(\Theta_{pred}(I_{t}^{v1}),\Theta_{obs}(I_{t}^{v2})).$ (3)
Here, $\Theta_{pred}$ and $\Theta_{obs}$ denote the predictor and observer networks of EgoAgent, respectively, $I_{t}^{v1}$ and $I_{t}^{v2}$ denote the two different views derived from the egocentric image $I_{t}$.
Overall objective function. Given an input token sequence containing $t$ time steps, we have the overall objective function $\mathcal{L}$ defined as:
$\small\mathcal{L}=\frac{1}{t}\sum_{k=0}^{t}{(\lambda_{rep}\mathcal{L}_{rep}+\lambda_{pred}\mathcal{L}_{pred}+\lambda_{act}\mathcal{L}_{act}),}$ (4)
where $\lambda_{rep}$, $\lambda_{pred}$, and $\lambda_{act}$ are the corresponding loss weights for representing the world, predicting future world states and actions, respectively.
3.3 Task-Specific Inference
After training, EgoAgent inherently acquires the abilities to solve the three tasks designed to mirror human cognition.
World state prediction. Given a sequence of egocentric video clips and human motion, EgoAgent is capable of predicting the future world state $S_{t+1}$ by $S_{t+1}=\Theta(I_{1},A_{1},...,I_{t},A_{t}).$ Once predicted, the future world image can then be retrieved by measuring similarity within the learned feature space, enabling applications that require foresight into upcoming visual states.
Human action prediction. From an egocentric perspective and informed by previous actions, EgoAgent can generate plausible future action $A_{t}$ represented as 3D human motion by $A_{t}=\Theta(I_{1},A_{1},...,A_{t-1},I_{t}).$ This predictive capability of human motion is essential for applications in humanoid robotics, virtual environments, and interactive gaming, where anticipating human-like movements is crucial.
Visual representation. EgoAgent learns world states as feature representations, allowing it to extract meaningful features from input images by $S=\Theta(I),$ where $\Theta$ denotes the EgoAgent model. These features can be directly applied to representation tasks such as image classification.
Table 1: Performance comparison across three tasks. (1) For world state prediction, we report the feature retrieval accuracy of the predicted next-frame features by the averaged Top1 accuracy and mAP on total timesteps $T_{s}=4$. (2) For human action prediction, we report the 3D motion prediction results on Ego-Exo4D by MPJPE (cm) and MPJVE (cm/s) with time gaps of 1/30 and 1/10 seconds. (3) For visual representation, we report the $k$-NN evaluation results on ImageNet-100 and ImageNet-1K by Top1 and Top5 accuracy (%). We evaluate with the officially released checkpoints of representation models trained on egocentric video datasets and train all motion prediction models from scratch. denotes the model trained by ourselves for a fair comparison.
Method Training Dataset World State Prediction 3D Human Motion Prediction Visual Representation
Ego-Exo4D Ego-Exo4D (30fps) Ego-Exo4D (10fps) ImageNet-100 ImageNet-1K
[…]
VIP [31] Ego4D 1.67 6.10 8.04 21.14 1.59 1.59
R3M [37] Ego4D 24.42 37.26 4.82 14.42 0.77 0.77
[…]
DoRA [44] WT 30.15 45.01 55.08 78.06 34.52 52.50
DoRA [44] WT+Ego-Exo4D 29.78 44.34 51.74 74.14 31.46 47.82
Diffusion Policy-C [13] Ego-Exo4D 28.07 206.96 27.95 114.75
Diffusion Policy-T [13] Ego-Exo4D 25.92 353.24 25.85 148.82
HumanMAC [11] Ego-Exo4D 19.21 94.22 17.68 77.43
siMLPe [22] Ego-Exo4D 13.33 81.94 12.20 60.65
EgoAgent-300M WT+Ego-Exo4D 43.01 58.06 12.92 82.18 11.89 59.96 55.14 76.56 34.65 51.42
EgoAgent-1B WT+Ego-Exo4D 46.43 61.96 12.51 81.45 11.65 58.99 56.48 78.12 35.84 53.03
4 Experiments
4.1 Experimental Setups
Datasets. We train EgoAgent on two egocentric datasets, WalkingTours (WT) [44] and Ego-Exo4D [20]. WT is a video-only dataset with around 1.5M high-resolution frames captured across cities worldwide. Ego-Exo4D has 221.26 hours of egocentric videos along with 376K manually labeled and 9.2M automatically generated 3D body poses. To eliminate projection differences between Ego-Exo4D and WT, we undistort the raw fisheye sensor data into a pinhole camera model. We also apply a 20-frame sliding window filter to extract continuous pose sequences, resulting in 1,410,119 clips of synchronized egocentric videos and 3D body poses. Based on the EgoPose split in Ego-Exo4D, we reserve the validation set of Ego-Exo4D-v2333The whole Ego-Exo4D dataset is divided into Ego-Exo4D-v1 and Ego-Exo4D-v2 on the official website (https://ego-exo4d-data.org). for evaluation and use the remaining clips for training, yielding 1,378,672 training clips and 31,447 evaluation clips.
Implementation Details. To alleviate the training burden, we sample one image frame every five frames while retaining all 3D body poses. Specifically, for each clip with 20 frames, we divide it into $T_{s}=4$ time steps and sample the first image frame from each step. This results in every time step consisting of one image frame paired with five frames of 3D body poses. We follow [44] to avoid cropping noisy positive pairs and adopt two global crops, six local crops, and the same augmentations consistent with DINO [10].
Following the configurations in [21], we train EgoAgent with two model sizes, 300M and 1B. We train EgoAgent on the processed dataset for 72,000 iterations with a linear warm-up of 1800 iterations. We set the base learning rate to $6\times 10^{-4}$ with a cosine decay scheduler. The loss weights are set as $\lambda_{rep}=2,\lambda_{pred}=1,\lambda_{act}=3$, respectively. All models are trained from scratch using FP16 to speed up. The training takes 25/60 hours in total with a batch size of 1920 on 32/48 NVIDIA A100 GPUs for EgoAgent-300M/EgoAgent-1B, respectively.
Figure 3: Retrieval results for egocentric world state prediction. Correctly retrieved images are marked with green boundaries.
4.2 World State Prediction
To evaluate the effectiveness of EgoAgent in predicting future states of the world, we perform an egocentric video prediction task with feature retrieval. Specifically, given the egocentric video frames and 3D human pose sequences at the current time step, we build a query set that stores the predicted features of the next frame and a gallery set that contains the features directly extracted from all video frames. At time step $t$, if the predicted world state $S_{t+1}^{\prime}$ in the query set can correctly retrieve the corresponding $S_{t+1}$ in the gallery set, we treat it as a successful prediction. For representation models, since there are no actions to condition the input image features, the predicted next state $S_{t+1}^{\prime}$ is exactly the state $S_{t}$ at time step $t$. To avoid retrieving this exact same state, when querying with input image $I_{t}$, we ignore the scene state $S_{t}$ in the gallery set. Following common practice in retrieval tasks, we use Top1 accuracy and mean average precision (mAP) as metrics.
As shown in Tab. 1, EgoAgent outperforms representation models by a large margin. Specifically, EgoAgent-300M improves the performance of DoRA by +12.86% in Top1 accuracy and +13.05% in mAP, indicating that it predicts the future states based not only on the semantic similarity of images. In contrast, representation models rely solely on semantic similarity, struggling with visually similar frames in egocentric videos. Furthermore, scaling up the model enhances performance, as EgoAgent-1B achieves further performance gains upon EgoAgent-300M with +3.42% Top1 accuracy and +3.90% mAP. Note that DoRA, when trained on our combined dataset, exhibits performance degradation in both world state prediction and visual representation tasks. We attribute this to its tracker’s inability to track abrupt object motions in Ego-Exo4D. For further analysis, please refer to Sec. 4.4. Fig. 3 visualizes the retrieval image by predicted features, where DoRA fails to retrieve the correct future states when humans have large movements, while EgoAgent successfully predicts the future state features that retrieve the real next-frame image.
Figure 4: Retrieval results for egocentric world state prediction from the same input observation with diverse human poses.
To further evaluate whether EgoAgent accurately models the causal relationship between actions and future world states, we conduct a qualitative analysis by conditioning next-state predictions on various poses with the same initial observing image, including dribbling, raising, shooting, and passing the ball. Fig. 4 shows the retrieved images, where all predicted features successfully retrieve images that accurately reflect the intended motion dynamics. This demonstrates EgoAgent’s ability to understand both the semantics of actions and their consequences in the egocentric world.
Figure 5: Visualizations of the 3D motion prediction tasks. After observing the egocentric frames, EgoAgent can generate accurate human skeletons even though most of the 3D body points are not visible in the input frame.
4.3 3D Human Motion Prediction
We compare EgoAgent with the state-of-the-art video-based motion generation model, Diffusion Policy [13] with convolution (C) and transformer (T) architectures, and unconditional motion prediction models [22, 11] using the EgoPose from Ego-Exo4D [20]. Each testing clip contains 5 frames of egocentric images and 20 frames of 3D human motion. We set the prediction target as the last 15 frames of 3D motion. The images and first 5 frames of 3D motion are adopted as input for Diffusion Policy and EgoAgent, while only the first 5 poses are used as input for unconditional motion prediction models. Following Ego-Exo4D [20], the mean per-joint position error (MPJPE) in centimeters (cm) and the mean per-joint velocity error (MPJVE) in centimeters per second (cm/s) are adopted as evaluation criteria.
Quantitatively, as shown in Tab. 1, EgoAgent achieves state-of-the-art performance with the lowest MPJPE and competitive MPJVE. Specifically, EgoAgent-300M improves siMLPe [22] by -0.41 and -0.31 MPJPE on the evaluation set of 30fps and 10 fps, respectively. When scaling up to 1B, the prediction errors are further reduced. These results demonstrate that EgoAgent can predict actions accurately when given egocentric observations and past actions.
We also qualitatively evaluate the motion prediction results. As illustrated in Fig. 5, EgoAgent produces relatively small errors compared to the ground truth. Compared to Diffusion Policy [13], EgoAgent achieves higher accuracy on non-visible joints. Specifically, in the early frames ($t_{1}$ and $t_{2}$), the missing right hand in the image leads to large prediction errors in Diffusion Policy, whereas EgoAgent accurately estimates hand positions by leveraging prior motion cues. Compared to HumanMAC [11], EgoAgent shows smaller accelerated errors. The predicted skeletons in HumanMAC gradually lean toward the floor ($t_{[3:5]}$) after several frames, whereas EgoAgent integrates visible body part information from egocentric observations to correct accelerated errors, leading to more stable motion predictions.
4.4 Visual Representation
We examine EgoAgent’s visual perception ability by two aspects: (1) image classification accuracy by $k$-NN on ImageNet-100 and ImageNet-1K [15], which is the golden standard for evaluating visual representations, and (2) egocentric manipulation success rate in an embodied simulator called TriFinger [48], which measures the transferability of learned visual representations to visual-motor control. Concretely, we freeze the model and extract features of images across two tasks, for image classification, we utilize a $k$-nearest neighbor classifier with $k=20$.
As shown in Tab. 1, EgoAgent achieves the best representation performance. Specifically, EgoAgent-1B outperforms the officially released DoRA by +1.40% and +1.32% Top1 accuracy on ImageNet-100 and ImageNet-1K, respectively, demonstrating that learning to predict world states and actions can facilitate the task of representing the world. For a fair comparison, we train DoRA on the same combined dataset (WT and Ego-Exo4D), which instead results in worse performance than training on WT alone. We suggest this is because, unlike WT videos, which feature stable motion, Ego-Exo4D includes fewer objects but more dynamic movements (objects frequently move out of sight). As a result, the tracker in DoRA struggles to maintain consistent object tracking, leading to degraded performance. In contrast, EgoAgent learns robust representation by modeling the temporal causality of state features and actions.
Table 2: Success rate (%) on the TriFinger [48] benchmark, where each model’s pretrained visual representation is fixed, and additional linear layers are trained as the policy network.
Methods Training Dataset Reach Cube Move Cube
[…]
DoRA [44] WT 82.40 48.13
DoRA [44] WT+Ego-Exo4D 81.55 45.41
EgoAgent-300M WT+Ego-Exo4D 82.61 54.21
EgoAgent-1B WT+Ego-Exo4D 85.72 57.66
Table 3: Ablation studies: removing one of the three tasks during training (b-d), training on a single task (e-g), and using pretrained image latents from VQGAN’s reconstruction-based latent space (h-i). All variants are trained with 14,400 iterations.
Method Training Task World State Prediction Human Motion Prediction Visual Representation
Ego-Exo4D Ego-Exo4D (@30fps) ImgNet-100 ImgNet-1K
[…]
(a) Baseline 37.77 53.72 14.49 88.61 41.64 22.28
(b) w/o $\mathcal{L}_{pred}$ 14.70 89.62 39.12 20.97
(c) w/o $\mathcal{L}_{act}$ 34.86 49.13 39.92 21.31
(d) w/o $\mathcal{L}_{rep}$ 25.90 40.17 14.49 88.82
(e) only $\mathcal{L}_{pred}$ 33.23 47.28
(f) only $\mathcal{L}_{act}$ 14.32 91.13
(g) only $\mathcal{L}_{rep}$ 40.80 21.56
(h) Pixel-level latent 20.62 37.12 13.57 84.86 1.00 0.10
(i) Pixel-level latent 15.63 23.85 16.25 100.49 31.20 13.22
The TriFinger benchmark [48] involves a three-finger robot performing two tasks (reach cube and move cube) from an egocentric view. We freeze the pretrained models and use a 3-layer MLP as the policy network, training each task with 100 demonstrations following [33]. As shown in Tab. 2, EgoAgent achieves the highest success rates on both tasks, outperforming the egocentric representation models DoRA [44] with increases of +3.32% and +3.9% respectively. This result shows that by incorporating human action prediction into the learning process, EgoAgent demonstrates the ability to learn more effective representations that benefit both image classification and embodied manipulation tasks. This highlights the potential of leveraging human-centric motion data to bridge the gap between visual understanding and actionable policy learning.
4.5 Ablation Study
To evaluate the effectiveness of EgoAgent’s specialized designs, we conduct several ablation studies using EgoAgent-300M with a short learning schedule of 14,400 iterations.
Joint learning of representation, prediction, and action. EgoAgent demonstrates promising performance across the three tasks. To understand how these tasks interact during training, we conduct leave-one-out experiments and one-task-only experiments, as shown in Tab. 3 (a-g). We observe several key findings. First, removing any one task degrades the performance of the other two, highlighting the complementary roles of these tasks in learning. Specifically, removing the representation loss causes the largest drop in state prediction (-11.87% in Top1 accuracy and -13.55% in mAP), while removing the prediction loss significantly weakens representation learning, reducing Top-1 accuracy by -2.52% on ImageNet-100 and -1.31% on ImageNet-1K. This suggests that state prediction and visual representation are closely interconnected. For motion prediction, excluding the prediction loss leads to the most significant decline, increasing MPJVE by 1.01. Second, training on individual tasks yields worse performances on each task than learning all three tasks together, except for MPJPE in motion prediction. We suspect that when trained only on motion prediction, the model focuses excessively on minimizing joint position errors. However, without context from the other tasks, it struggles to maintain smooth, consistent motion across frames, resulting in a higher MPJVE of 2.52.
Analysis of task dependencies. We also identify interesting dependencies between the three tasks. The key interactions are that actions rely on the understanding (or representation) of the current world state, while next-state predictions depend on both the current state understanding and actions taken. Specifically, (1) Representation serves as the foundation for both prediction and action as it captures semantic information to model the current world state. (2) Prediction and action jointly provide effective supervision for representation by modeling world dynamics and predicting actions, improving representation by 1.16% in Top1 accuracy. However, training on prediction or action alone cannot enhance representation, as actions without accurate next-state predictions or next-state predictions conditioned on incorrect actions may provide noisy supervision. (3) Representation and action guide the model to integrate learned representations with action dynamics, leading to more accurate predictions of future states. Meanwhile, representation and prediction work together to help the model understand how actions influence state transitions, refining its ability to generate actions.
Training in semantic feature space. To evaluate the effectiveness of training agent models with a high-level semantic feature space, we use a pretrained VQGAN [17] as the image tokenizer. The results are shown in Tab. 3 (a,h,i). Training EgoAgent in VQGAN’s pretrained pixel-level latent space results in significant performance drops in world state prediction and nearly nullifies the ability for visual representation. While motion prediction improves, likely due to the unified task formulation, the overall performance degradation makes this latent space unsuitable. Forcing representation learning with this pixel-level latent space not only yields modest representation performance but also significantly weakens the other two tasks compared to learning in continuous semantic features. This suggests that high-level semantics are more effective than low-level details for understanding world states and distinguishing different contents in adjacent timesteps.
5 Conclusion and Future Work
In this paper, we presented EgoAgent, a joint predictive agent model that simultaneously learns to represent world observations, predict future world states, and generate 3D human motions in egocentric worlds. EgoAgent integrates these three abilities within a joint embedding-action-prediction architecture. Through extensive experiments, we demonstrated that EgoAgent outperforms existing state-of-the-art methods, with the joint learning of perception, prediction, and action tasks mutually reinforcing each other, leading to improved performance across all three capabilities. Future work will explore incorporating finer-grained hand representations for precise tasks such as object manipulation and integrating long-term memory mechanisms for tasks requiring extended temporal dependencies.